Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Where did I go wrong?

Expert replies
by winniethepooh » Wed Jun 15, 2011 9:57 pm
Q.In a bag of marbles, there are 3 red, 2 white, and 5 blue marbles. If Maddy takes 2 marbles out of the bag, what is the probability that he will have one white and one blue marble?


What I have done is calculated probability of taking out 1 white marble that is 2/10 and multiplied that with the probability of taking out the blue marble that is 5/9..Multiplying 2/10 and 5/9 I get 1/9, which is not the answer!!
Can I know, whether is my approach wrong or did I go wrong somewhere?
Join the discussion
Source: — Problem Solving |

by Geva@EconomistGMAT » Wed Jun 15, 2011 10:06 pm
winniethepooh wrote:Q.In a bag of marbles, there are 3 red, 2 white, and 5 blue marbles. If Maddy takes 2 marbles out of the bag, what is the probability that he will have one white and one blue marble?


What I have done is calculated probability of taking out 1 white marble that is 2/10 and multiplied that with the probability of taking out the blue marble that is 5/9..Multiplying 2/10 and 5/9 I get 1/9, which is not the answer!!
Can I know, whether is my approach wrong or did I go wrong somewhere?
I'm guessing that the right answer is twice that, or 2/9: your calculation assumes that Maddy draws white first, blue second, but the question asks for "one white and one blue" - which doesn't really imply order. Thus, you need to bring both scenarios into account:

white, then blue: 2/10 * 5/9 = 1/9
Blue, then white: 5/10 * 2/9 = 1/9

total probability = 1/9+1/9=2/9.
Geva
Senior Instructor
Master GMAT
1-888-780-GMAT
https://www.mastergmat.com
Join the discussion

by Frankenstein » Wed Jun 15, 2011 10:08 pm
Hi,
You have taken white marble first and blue marble next. He has actually taken both the balls at the same. So, you have to add the case of picking blue marble first and then white marble as well.
So, it will be (2/10)(5/9) + (5/10)(2/9) = 2/9
(or)
Selecting 1 white and 1 blue can be done in 2C1*5C1
Number of ways of selecting 2 balls from 10 is 10C2.
So, probability is 2C1*5C1/10C2 = 2/9
Cheers!

Things are not what they appear to be... nor are they otherwise
Join the discussion